OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 5x. What is the value of (f o g)(3)? A. 3x + 8 B. -22 C. 29 D. 17 is it B?

OpenStudy (anonymous):

(f o g)(3) means "f of g of 3" which is the same as f(g(3)) plug 3 into g, and then plug that into f. g(3) = 5*3 = 15 f(g(3)) = f(15) = -2(15)+8 = -30+8 = -22 So, yes. Good work =)